What is the volumn in gallons of a tank with a 6' raduius and 3'deep?
Found 2 solutions by Alan3354, checkley77:
Answer by Alan3354(69443)

Vol = pi*r^2*h
V = pi*36*3 cubic ft = 108pi cub ft
V =~ 2538 gallons (using 231 cubic inches per gallon)

Answer by checkley77(12844)

V=BH
V=2PIR*H
V=2*3.14*6*3
V=113.04 fT^3
.35 FT^3=2.64 GAL.
(113.04/.35)*2.64
645.94*2.64
1,705.3 GAL. ANS.